Presented below is information related to Starr Company.

1. Net Income [including a discontinued operations gain (net of tax) of $69,000] $229,500
2. Capital Structure
a. Cumulative 5% preferred stock, $100 par, 6,300 shares issued and outstanding $630,000
b. $10 par common stock, 74,000 shares outstanding on January 1. On April 1, 40,000 shares were issued for cash. On October 1, 16,000 shares were purchased and retired. $1,000,000
c. On January 2 of the current year, Starr purchased Oslo Corporation. One of the terms of the purchase was that if Oslo net income for the following year is $237,000 or more, 50,000 additional shares would be issued to Oslo stockholders next year. Oslo’s net income for the current year was $2,600,000.
3. Other Information
a. Average market price per share of common stock during entire year $30
b.

Income tax rate 30%

1. Compute weighted average shares outstanding.

Weighted average shares outstanding

2. Compute earnings per share for the current year. (Round answers to 2 decimal places, e.g. 52.75.)

Basic earnings per share $
Diluted earnings per share $   
Answer #1

Solution:

1)Calculation of weighted average share outstanding is given below

Number of shares outstanding(A) Period for which the shares remained outstanding B Total months(C) Weighted shares(A*B/C)
1st January 74,000 3 (1st January to 31st March) 12 18,500
1st April 114,000(74,000 + 40,000) 6 (1st April to 30th September) 12 57,000
1st October 98,000 3 (1st October to 31st December) 12 24,500
Total weighted average shares outstanding 100,000 shares

Weighted-Average Shares Outstanding = 100,000 shares

2)

To determine the EPS, we first need to calculate the net income available to common stock-holders. The calculation of net income is given in the following table:

Income Before Extraordinary Item ($229,500 -69,000) $160,500
Less: Preferred Dividends (630,000 * 5%) $31,500
Income Available to Common Stockholders before Extraordinary Item 129,000
Add: Extraordinary Gain (Net of Tax) 69,000
Income Available to Common Stockholders $198,000

The formula for calculating Basic EPS and Diluted EPS are given below:

Basic EPS = Income Available to Common Stockholders/Weighted Average Shares Outstanding

Diluted EPS = Income Available to Common Stockholders/Weighted Average Shares Outstanding + Dilutive Additional Shares

Basic EPS = 198,000 / 100,000 = $1.98 per share

Dilutive EPS = 198,000 /(100,000 + 50,000) = $1.32 per share